# Branch length statistics in phylogenetic trees under constant-rate birth-death dynamics

### v3

@@ I Introduction
In Stadler and Steel (2012); Mooers et al. (2012) Stadler, Steel and collaborators focus on a particular aspect of tree statistics, namely the distribution of branch lengths in calendar time. They calculate the probability density function for the length of the so-called pendant branches (the lowest branches leading to the leaves of the tree) for trees reconstructed from a given number of leaves. For the particular case of a Yule process (a birth-death process with the rate of death set to zero), they also give the probability density function for the length of interior branches. In Paradis (2016) Paradis introduces an approach to calculate branch length distributions of both pendant and interior branches at a particular time since the start of the birth-death process.
@@
In this paper, we revisit the problem of the distribution of branch lengths with a particular focus on the trees reconstructed from a finite fraction of extant individuals sampled at the time of observation. We distinguish two distinct statistical scenarios: In the first scenario (scenario i)), the time interval between the start of the birth-death process with a single individual and the time of observation (the end of the birth-death process) is known, but the number of extant individuals at the end of the process is a fluctuating random variable. In the second scenario (scenario ii)), one conditions on the number of individuals at the end of the process, but the time interval between the start and end of the process is unknown.
@@
In Section II, we recapitulate the relevant properties of a constant-rate birth-death process. In Sections III-V we use the approach introduced by Paradis Paradis (2016) to systematically write down concrete expressions for branch length distributions for different types of branches under the first scenario: In Section III we focus on pendant branches, in Section IV on branches one and two levels above a pendant branch, and in Section V on generic branches in the interior of a phylogenetic tree. Sections III and V coincide with the work of Paradis in Paradis (2016); they are included here for completeness and because Paradis (2016) does not give concrete expressions for the branch length distributions. Section VI gives the key results of this paper; we consider branch lengths arising in trees reconstructed from only a finite fraction of the extant population. We find that sampling affects different parts of the reconstructed tree differently: while pendant branches become longer and grow without limit as the sampling probability is taken to zero, the distribution of interior branch lengths quickly reaches an asymptotic limit as the sampling probability decreases. In Section VII we turn to the second scenario (conditioning on the number of extant individuals at the time of observation) and introduce a novel approach based on tracking the birth-death process first backwards in time from a given number of individuals at the time of observation to the emergence of a particular branch and then again forwards again to the time of observation. For pendant branches, we recover the results of Stadler, Steel and collaborators Stadler and Steel (2012); Mooers et al. (2012) as expected and generalize these results to interior branches.
@@ II Dynamics of a birth-death process
A powerful alternative to the master equation (1) is a master equation for the probability that a population that started with a single individual at time zero has population size one at time $t$ Maddison et al. (2007); Stadler (2010). A generalization of this approach is the probability that in a population that started with a single individual at time zero, at time $t$ there is one and only one individual that has surviving offspring at a later time $T$. This probability turns out to be a key tool to calculate the distribution of branch lengths.
@@
Specifically, $\tilde{p}^{(T)}_{1}(t)$ denotes the probability that at time $t$ a clade which started at time zero contains one individual with open future fate, and all other individuals alive at time $t$ have no extant descendants at time $T$ (including themselves). The master equation for $\tilde{p}^{(T)}_{1}(t)$ is
$\partial_{t}\tilde{p}^{(T)}_{1}(t)=-\mu\tilde{p}^{(T)}_{1}(t)-\lambda\left(1-2p_{0}(T-t)\right)\tilde{p}^{(T)}_{1}(t)\ ,$ (6)
@@
where the first term accounts for the decay of probability due to that individual’s death, and the second term describes the probability that any additional lineage arising in a birth event fails to become extinct by the time of observation. The factor of two comes from the two offspring in such a birth event which can play this role. This is a master equation in a single probability only; the population aspect of this dynamics resides compactly in the probability $p_{0}(T-t)$ that a particular clade dies out in a time interval of length $T-t$ given by (2).
@@
The master equation (6) is a first-order linear differential equation solved by the integrating factor $-\int_{0}^{t}\operatorname{\mathrm{d}\!}t^{\prime}\left[\mu+\lambda\left(1-2p_{0}(T-t^{\prime})\right)\right]$. Integrating to $t=T$ yields the probability of population size one at time $T$ (3). Integrating $t$ only to some intermediate time $T-\tau_{e}$ gives the probability $\tilde{p}^{(T)}_{1}(T-\tau_{e})$. (The name $\tau_{e}$ is chosen as this will later mark the end of an interior branch.)
@@
It turns out to be convenient for the analysis of branch lengths to consider times relative to the time of observation $T$. We denote the length of the interval between the intermediate time $t$ and the observation time $T$ by $\tau_{e}=T-t$. We further define $\tau=T-\tau_{e}$ as the length of the interval between $\tau_{e}$ and the time of the clade’s origin. Specifically, we denote as $p_{1}^{(\tau_{e})}(\tau)=\tilde{p}^{(T)}_{1}(t=T-\tau_{e})$ the probability that a clade started at time $\tau_{e}+\tau$ in the past, and contains at time $\tau_{e}$ in the past only a single individual that could have extant descendants at the time of observation. Again, setting $\tau_{e}=0$ results in $p_{1}^{(\tau_{e}=0)}(\tau)\equiv p_{1}(\tau)$. For the case of $\tau_{e}>0$ the integrating factor becomes $-\int_{0}^{T-\tau_{e}}\operatorname{\mathrm{d}\!}t^{\prime}\left[\mu+\lambda\left(1-2p_{0}(T-t^{\prime})\right)\right]=-\int_{0}^{\tau}\operatorname{\mathrm{d}\!}\tau^{\prime}\left[\mu+\lambda\left(1-2p_{0}(\tau_{e}+\tau^{\prime})\right)\right]=\ln\left(\frac{(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{e}})^{2}\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}}{(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)(\tau_{e}+\tau)})^{2}}\right)$ and we obtain
@@
Figure 1: Times of events on a birth-death tree. Sketch of the situation characterized by $p_{1}^{(\tau_{e})}(\tau)=\tilde{p}_{1}^{(T)}(t)$. From the perspective of a clock starting at the origin of a clade, $p_{1}^{(\tau_{e})}(\tau)=\tilde{p}_{1}^{(T)}(t)$ describes the probability that out of all individuals at time $t$, only one will have extant descendants (depicted in blue) at time $T>t$ when the population is observed. The red lines show lineages which die out before $T$. For convenience, the time of observation is used as the origin of time $\tau$, with time increasing as we look into the evolutionary past. $\tau_{e}$ denotes time point beyond only a single individual has a clade extant at the time of observation, and $\tau$ denotes the time the population has had to grow before $\tau_{e}$. In the corresponding reconstructed phylogenetic tree, there are only the black and blue lineages.
@@ III Branch length distribution: pendant branches
A pendant branch (a branch terminating at the time of observation) of length $\tau$ was created in a birth event at time $\tau$ , where (i) one of the two offspring produces a clade consisting of a single extant member at the time of observation, and (ii) the other offspring produces a clade that did not die out by the time of observation. Figure 2A shows a simple example. Conditions (i) and (ii) are statistically independent events, whose probabilities can be expressed in terms of the quantities computed above. This leaves the rate at which birth events occur per small time interval, which is given by the birth rate $\lambda$ times the population size at time $\tau$. Since the population size is unknown, we sum over all possible population sizes $n$ and their probabilities, yielding that the probability of a birth event at time $\tau$ as the expectation value of the population size multiplied with the birth rate $\lambda$. Plugging $t=T-\tau$ in equation (5) gives the expected population size at time $\tau$ before sampling is proportional to $\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}$.
The relative number of pendant branches of length $\tau$ per small time interval is thus given by
$\lambda\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}\left[p_{1}(\tau)p_{>1}(\tau)+p_{>1}(\tau)p_{1}(\tau)+2p_{1}(\tau)^{2}\right]$ (8)
$=2\lambda\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}p_{1}(\tau)\left(1-p_{0}(\tau)\right).$ (9)
The factor of two at the end of the first term accounts for the emergence of two branches of length $\tau$ when both offspring only have a single surviving descendant. $p_{>1}(\tau)=\sum_{n=2}^{\infty}p_{n}(\tau)$ denotes the probability that a clade founded at time $\tau$ has more than one extant member at the time of observation. Using equations (2) and (3), one obtains the probability density function of the lengths $\tau$ of pendant branches as
@@
For a finite time since the start of the birth-death process, $\tau$ has a finite support and hence $P_{\text{pend}}(\tau)$ cuts off to zero when $\tau$ is larger than the time $T$ since the start of the birth-death process. The normalizing factor of (10) is then found to be $(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)T})^{2}/(\lambda(\lambda-(\lambda-2\mu)\operatorname{\mathrm{e}}^{-2(\lambda-\mu)T}-2\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)T}))$. We focus on the particularly simple case of an asymptotically long running time $T\to\infty$. In this case, the cutoff moves to infinity and we find the normalizing factor in (10) to be one, making (10) a proper probability density function. Since the normalizing factor converges exponentially towards one, the asymptotic distribution is also a valid approximation for populations of finite age $T$ as long as $(\lambda-\mu)T\gg 1$. We will focus on this asymptotic case also when treating branches in the tree’s interior.
Curiously, distribution (10) coincides with the distribution of pendant branches conditioned on the population size $N$ (statistical scenario ii)), derived by Stadler and Steel Stadler and Steel (2012) as well as Mooers et al. Mooers et al. (2012). For details, see Appendix .1.
@@ IV Branch length distribution: interior branches above pendant branches
This approach to pendant branches can be extended to the interior branches of a phylogenetic tree. We proceed to go up the phylogenetic tree and first consider branches one level up from pendant branches. We call branches that end in an internal node that is the beginning of a pendant branch level-two branches. Correspondingly, pendant branches are level-one branches. Figure 2B shows the situation: a pair of successive level-two and level-one branches of lengths $\tau_{2}$ and $\tau_{1}$, respectively, arose in a birth event at time $\tau_{2}+\tau_{1}$ in the past, where one of the two offspring has extant descendants (d-f), and the other has no extant descendants apart from the offspring of individuals arising in another birth event at time $\tau_{1}$, which gives rise to two offspring. Again, one of these has at least one extant descendant (b-c), and the other has exactly one (a), establishing the pendant branch of length $\tau_{1}$, see Fig. 2B.
ARefer to caption                    BRefer to caption
CRefer to caption
@@
Since the probability of a birth event at time $\tau_{1}+\tau_{2}$ was $\lambda\operatorname{\mathrm{e}}^{-(\lambda-\mu)(\tau_{1}+\tau_{2})}$ (up to a normalizing prefactor, see below), the relative number of level-two and level-one branches of lengths $\tau_{2}$ and $\tau_{1}$ per small time intervals is thus
@@
Here $p^{(\tau_{1})}_{1}(\tau_{2})$ denotes the probability that all birth events taking place along the level-two branch of length $\tau_{2}$ lead to additional lineages that die out by the time the population is observed. More precisely, it is the probability that a given individual alive at time $\tau_{1}+\tau_{2}$ has exactly one descendant alive at time $\tau_{1}$ that will have extant descendants. It is given by equation (7).
Equation (11) gives the joint probability density function of branch lengths of pendant and level-two branches up to a normalization factor, which again turns out to be one. Figure 2D compares this result to numerical simulations. A population with birth rate $\lambda=1$ and death rate $\mu=0.75$ was grown from a single individual until time $T=25$ (where the population contained $4460$ individuals), and the phylogenetic tree was reconstructed. The lengths of level-two branches were collected in a histogram and compared to the joint probability $P_{\text{level}}(\tau_{1},\tau_{2})$ given by (11) marginalized with respect to $\tau_{1}$.

### v4

@@ I Introduction
In Stadler and Steel (2012); Mooers et al. (2012) Stadler, Steel and collaborators focus on a particular aspect of tree statistics, namely the distribution of branch lengths in calendar time. They calculate the probability density function for the length of the so-called pendant branches (the lowest branches leading to the leaves of the tree) for trees reconstructed from a given number of leaves. For the particular case of a Yule process (a birth-death process with the rate of death set to zero), they also give the probability density function for the length of interior branches. In Paradis (2016), Paradis introduces an approach to calculate branch length distributions of both pendant and interior branches at a particular time since the start of the birth-death process.
@@
In this paper, we revisit the problem of the distribution of branch lengths with a particular focus on the trees reconstructed from a finite fraction of extant individuals sampled at the time of observation. We find that when the sampled fraction of the population is lowered, pendant branches increase in length while interior branches converge to an asymptotic distribution which is independent of the sampling fraction. In our analysis, we distinguish two distinct statistical scenarios: In the first scenario (scenario i)), the time interval between the start of the birth-death process with a single individual and the time of observation (the end of the birth-death process) is known, but the number of extant individuals at the end of the process is a fluctuating random variable. In the second scenario (scenario ii)), one conditions on the number of individuals at the end of the process, but the time interval between the start and end of the process is unknown.
@@
In Section II, we recapitulate the relevant properties of a constant-rate birth-death process. In Sections III-V we use the approach introduced by Paradis Paradis (2016) to systematically write down concrete expressions for branch length distributions for different types of branches under the first scenario: In Section III we focus on pendant branches, in Section IV on branches one and two levels above a pendant branch, and in Section V on generic branches in the interior of a phylogenetic tree. Sections III and V coincide with the work of Paradis in Paradis (2016); they are included here for completeness and because Paradis (2016) does not give concrete expressions for the branch length distributions. Section VI gives the key results of this paper; we consider branch lengths arising in trees reconstructed from only a finite fraction of the extant population. We find that sampling affects different parts of the reconstructed tree differently: while pendant branches become longer and grow without limit as the sampling probability is taken to zero, the distribution of interior branch lengths quickly reaches an asymptotic limit as the sampling probability decreases. In Section VII we turn to the second scenario (conditioning on the number of extant individuals at the time of observation) and introduce a novel approach based on tracking the birth-death process first backwards in time from a given number of individuals at the time of observation to the emergence of a particular branch and then again forwards again to the time of observation. For pendant branches, we recover the results of Stadler, Steel and collaborators Stadler and Steel (2012); Mooers et al. (2012) and generalize these results to interior branches.
@@ II Dynamics of a birth-death process
A powerful alternative to the master equation (1) is a master equation for the probability that a population that started with a single individual at time zero has population size one at time $t$ Maddison et al. (2007); Stadler (2010). A generalization of this approach is the probability that in a clade that started with a single individual at time $t_{0}$, at time $t_{e}$ there is one and only one individual that has surviving offspring at a later time $T$. This probability turns out to be a key tool to calculate the distribution of branch lengths.
@@
Specifically, $\tilde{p}^{(t_{0},T)}_{1}(t_{e})$ denotes the probability that at time $t_{e}$ a clade which started at time $t_{0}$ contains one individual with open future fate, and all other individuals alive at time $t_{e}>t_{0}$ have no extant descendants at time $T$ (including themselves). This definition appears cumbersome at first, but is crucial to characterize a branch; all lineages that split off from the branch in bifurcations between its start $t_{0}$ and end $t_{e}$ must die out by the time of observation, otherwise the branch ends before $t_{e}$. The master equation for $\tilde{p}^{({t_{0},}T)}_{1}({t_{e}})$ is
$\partial_{{t_{e}}}\tilde{p}^{({t_{0},}T)}_{1}({t_{e}})=-\mu\tilde{p}^{({t_{0},}T)}_{1}({t_{e}})-\lambda\left(1-2p_{0}(T-{t_{e}})\right)\tilde{p}^{({t_{0},}T)}_{1}({t_{e}})\ ,$ (6)
@@
where the first term accounts for the decay of probability due to that individual’s death, and the second term describes the probability that any additional lineage arising in a birth event fails to become extinct by the time of observation. The factor of two comes from the two offspring in such a birth event which can play this role. This is a master equation in a single probability only; the population aspect of this dynamics resides compactly in the probability $p_{0}(T-{t_{e}})$ that a particular clade dies out in a time interval of length $T-{t_{e}}$ given by (2).
@@
The master equation (6) is a first-order linear differential equation solved by an exponential function with the integrating factor $\int_{t_{0}}^{{t_{e}}}\operatorname{\mathrm{d}\!}t^{\prime}\left[\mu+\lambda\left(1-2p_{0}(T-t^{\prime})\right)\right]$ as its argument. Integrating to ${t_{e}}=T$ yields the probability of population size one at time $T$ (equation (3) when $t_{0}=0$). Integrating $t_{e}$} only to some intermediate time $T-\tau_{e}$ gives the probability $\tilde{p}^{({t_{0},}T)}_{1}(T-\tau_{e})$. (The name $\tau_{e}$ is chosen as this will later mark the end of an interior branch.)
@@
It turns out to be convenient for the analysis of branch lengths to consider times relative to the time of observation $T$. We denote the length of the interval between the intermediate time $t_{e}$} and the observation time $T$ by $\tau_{e}=T-{t_{e}}$. We further define $\tau=T{-t_{0}}-\tau_{e}$ as the length of the interval between $\tau_{e}$ and the time of the clade’s origin. Specifically, we denote as $p_{1}^{(\tau_{e})}(\tau)=\tilde{p}^{({t_{0},}T)}_{1}({t_{e}}=T-\tau_{e})$ the probability that a clade started at time $\tau_{e}+\tau$ in the past, and contains at time $\tau_{e}$ in the past only a single individual that could have extant descendants at the time of observation while all additional lineages die out before observation (see explanation above for $\tilde{p}^{({t_{0},}T)}_{1}({t_{e}})$). Again, setting $\tau_{e}=0$ results in $p_{1}^{(\tau_{e}=0)}(\tau)\equiv p_{1}(\tau)$. For the case of $\tau_{e}>0$ the integrating factor becomes $-\int_{t_{0}}^{T-\tau_{e}}\operatorname{\mathrm{d}\!}t^{\prime}\left[\mu+\lambda\left(1-2p_{0}(T-t^{\prime})\right)\right]=-\int_{0}^{\tau}\operatorname{\mathrm{d}\!}\tau^{\prime}\left[\mu+\lambda\left(1-2p_{0}(\tau_{e}+\tau^{\prime})\right)\right]=\ln\left(\frac{(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{e}})^{2}\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}}{(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)(\tau_{e}+\tau)})^{2}}\right)$ and we obtain
@@
Figure 1: Times on a birth-death tree. At time $t_{0}$ (with time running forwards), a clade is founded by a single individual; birth and death events then change the number of individuals in the clade over time. We focus on the probability $\tilde{p}_{1}^{({t_{0},}T)}(t_{e})$ that the clade at time $t_{e}$ has the following property: all individuals and their descendants will die out by the time of observation $T$ (lineages shown in red), except for one particular individual (black lineage) whose fate is unspecified. Here this individual will later give rise to two extant individuals shown in blue. In the corresponding reconstructed phylogenetic tree, there are only the black and blue lineages. For convenience, we define $\tau_{e}=T-t_{e}$ (which runs backwards, thus increasing as we go into the evolutionary past), and $\tau$ as the interval between the origin of the clade and $\tau_{e}$. Using this notation, the probability defined above is written as $p_{1}^{(\tau_{e})}(\tau)=\tilde{p}^{(t_{0},T)}_{1}(t_{e}=T-\tau_{e})$, see text.
@@ III Branch length distribution: pendant branches
A pendant branch (a branch terminating at the time of observation) of length $\tau$ is defined by a birth event at time $\tau$, where (i) one of the two offspring produces a clade consisting of a single extant member at the time of observation, and (ii) the other offspring produces a clade that did not die out by the time of observation. Figure 2A shows a simple example. Conditions (i) and (ii) are statistically independent events, whose probabilities can be expressed in terms of the quantities computed above. This leaves the rate at which birth events occur per small time interval, which is given by the birth rate $\lambda$ times the population size at time $\tau$. The probability of two or more birth events occurring in the same time interval vanishes quadratically with the small time interval and can be neglected (compared to the probability of a single birth event). The population size is a random variable, and we sum over all possible population sizes $n$ and their probabilities, yielding the rate of a birth event at time $\tau$ as the expectation value of the population size multiplied with the birth rate $\lambda$. Plugging $t=T-\tau$ in equation (5) gives the expected population size at time $\tau$ before sampling is $\operatorname{\mathrm{e}}^{(\lambda-\mu)(T-\tau)}$.
The expected number of pendant branches of length $\tau$ per small time interval is thus given by
$\lambda\operatorname{\mathrm{e}}^{(\lambda-\mu){(T-\tau)}}\left[p_{1}(\tau)p_{>1}(\tau)+p_{>1}(\tau)p_{1}(\tau)+2p_{1}(\tau)^{2}\right]$ (8)
$=2\lambda\operatorname{\mathrm{e}}^{(\lambda-\mu){(T-\tau)}}p_{1}(\tau)\left(1-p_{0}(\tau)\right).$ (9)
The factor of two at the end of the first term accounts for the emergence of two branches of length $\tau$ when both offspring only have a single surviving descendant. $p_{>1}(\tau)=\sum_{n=2}^{\infty}p_{n}(\tau)$ denotes the probability that a clade founded at time $\tau$ has more than one extant member at the time of observation. The corresponding relative frequencies of branch lengths give the probability distribution of $\tau$ up to a prefactor (to be determined by normalization). Using equations (2) and (3), we obtain the probability density function of branch lengths up to a normalizing factor (see below)
@@
This is the probability density function (probability per small time interval) of the lengths of pendant branches for the case of asymptotically long running times $T\to\infty$, where the normalizing factor turns out to be one for the interval $0<\tau<T=\infty$. For a finite time since the start of the birth-death process, $\tau$ has a finite support and hence $P_{\text{pend}}(\tau)$ cuts off to zero when $\tau$ is larger than the time $T$ since the start of the birth-death process. The normalizing factor of (10) is then found to be $(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)T})^{2}/(\lambda(\lambda-(\lambda-2\mu)\operatorname{\mathrm{e}}^{-2(\lambda-\mu)T}-2\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)T}))$. Since the normalizing factor converges exponentially towards one as $T$ increases, the asymptotic distribution (10) is also a valid approximation for populations of finite age $T$ in the regime $(\lambda-\mu)T\gg 1$.
@@ IV Branch length distribution: interior branches above pendant branches
This approach to pendant branches can be extended to the interior branches of a phylogenetic tree. We proceed to go up the phylogenetic tree and first consider branches one level up from pendant branches. We call branches that end in an internal node that is the beginning of a pendant branch level-two branches. Correspondingly, pendant branches are level-one branches. Figure 2B shows the situation: a pair of successive level-two and level-one branches of lengths $\tau_{2}$ and $\tau_{1}$, respectively, arose in a birth event at time $\tau_{2}+\tau_{1}$ (with time running backwards as described above), where one of the two offspring has extant descendants (d-f), and the other has no extant descendants apart from the offspring of individuals arising in another birth event at time $\tau_{1}$, which gives rise to two offspring. Again, one of these has at least one extant descendant (b-c), and the other has exactly one (a), establishing the pendant branch of length $\tau_{1}$, see Fig. 2B.
ARefer to caption  BRefer to caption  CRefer to caption
@@
We proceed analogously to pendant branches: A a pair of successive level-two and level-one branches is created by a birth event at time $\tau_{1}+\tau_{2}$, the rate of which we know. This rate is multiplied by the probabilities that a level-two and level-one branch are created (as specified above). Since the probability of a birth event at time $\tau_{1}+\tau_{2}$ before observation was $\lambda\operatorname{\mathrm{e}}^{-(\lambda-\mu)(\tau_{1}+\tau_{2})}$ (up to a normalizing prefactor, see below, and per small time interval), the relative number of level-two and level-one branches of lengths $\tau_{2}$ and $\tau_{1}$ per small time intervals is thus
@@
$p^{(\tau_{1})}_{1}(\tau_{2})$ is given by (7) and denotes the probability that all birth events taking place along the level-two branch of length $\tau_{2}$ lead to additional lineages that die out by the time the population is observed. More precisely, it is the probability that a given individual alive at time $\tau_{1}+\tau_{2}$ has exactly one descendant alive at time $\tau_{1}$ that will have one or more extant descendants.
Equation (11) gives the joint probability density function of branch lengths of pendant and level-two branches up to a normalization factor, which again turns out to be one in the limit of large times $T$. Figure 2D compares this result to numerical simulations. A population with birth rate $\lambda=1$ and death rate $\mu=0.75$ was grown from a single individual until time $T=25$ (where the population contained $4460$ individuals), and the phylogenetic tree was reconstructed. The lengths of level-two branches were collected in a histogram and compared to the joint probability $P_{\text{level}}(\tau_{1},\tau_{2})$ given by (11) marginalized with respect to $\tau_{1}$.